MLC= 3+4L

calculate the cost of employing the first eight labourers ?

please show workings


1
Expert's answer
2021-06-01T11:27:32-0400

Solution:

MLC = 3 + 4L

L = Unit of labor

The cost of employing one labourer = 3 + 4(1) = 3 + 4 = 7

The cost of employing one labourer = 3 + 4(2) = 3 + 8 = 11

The cost of employing one labourer = 3 + 4(3) = 3 + 12 = 15

The cost of employing one labourer = 3 + 4(4) = 3 + 16 = 19

The cost of employing one labourer = 3 + 4(5) = 3 + 20 = 23

The cost of employing one labourer = 3 + 4(6) = 3 + 24 = 27

The cost of employing one labourer = 3 + 4(7) = 3 + 28 = 31

The cost of employing one labourer = 3 + 4(8) = 3 + 32 = 35

Total cost of employing the first 8 labourers = 7 + 11 + 15 + 19 + 23 + 27 + 31 + 35 = 149


The cost of employing the first eight laborers = 149